Copper(I) ions in aqueous solution react with NH3(aq)NH3(aq) according to

Cu+(aq)+2NH3(aq)⟶Cu(NH3)+2(aq)?f=6.3×1010Cu+(aq)+2NH3(aq)⟶Cu(NH3)2+(aq) Kf=6.3×1010

Calculate the solubility (in g·L−1) of CuBr(s)CuBr(s) (?sp=6.3×10−9) in 0.630.63 M NH3(aq)NH3(aq).
Answer #1

first to find k.

using kf and Ksp equation.

then to draw ICE table to find the molar solublity.

Gut + 2NWE → Cu(NU, K+ = 6138,010 key = {cul M6) = 63.XID— ) [cut JNK ₂] to find solublity of Cu Bo in 0.63M. NU Ksp [cat] [. molar solublity is 0.3148 M.

molar mass of CuBr = 143.45 g/mol

in g/L = 0.3148× 143.45 = 45.158 g/L